A study is designed to test H0: p = .5 against Ha: p > .5. A sample size of 100 and a significance level of 0.05 will be used.

If p = .6, find P(Type 2 error).
Answer #1
hypothesized proportion po= 0.5
true proportion pa= 0.6
sample size n= 100
standard error of po=√(po*(1-po)/n)= 0.0500
standard error of pa=√(pa*(1-pa)/n)= 0.0490
0.05 level and right tailed test critical value Zα=1.645
rejection reg:p̂ >=po+Zα*σpo or p̂ >0.5+1.645*0.05 = 0.5823
type II error=β=P( p̂ <0.58225|p=0.6)=P(Z<(0.58225-0.6)/0.049)=P(Z<-0.36)=0.3594

(please try 03585 if above comes wrong and reply)